# Semiregular abstract polyhedra with trivial facet stabilizer

### v1

@@ 9 Alternating semiregular polyhedra with trivial facet stabilizer
In Section 8.1 we have discussed how polyhedra with symmetry type graph in family 1 can be obtained by applying the voltage operation illustrated in Figure 8 to a regular polytope of rank $n$. If instead we apply this same operation to a chiral polytope $\mathcal{P}$ with automorphism group $\Gamma(\mathcal{P})=\langle\tau_{1},\tau_{2},\ldots,\tau_{n-1}\rangle$, where once again $\Phi\tau_{i}=r_{i}r_{0}\Phi$ for all $i$ and some fixed base flag $\Phi$, then [8, Theorem 6.1] tells us that the result is the derived graph of the voltage graph in Figure 16. This graph is obtained by taking two $n$-gons and then connecting corresponding edges with a ‘‘cross’’ of edges of color 2. That is, the left flag of the $i$-th edge of one copy is 2-adjacent to the right flag of the $i$-th edge of the second copy and viceversa. The darts of color $2$ that start on the $i$-th edge first copy have voltage $\tau_{i}$, and the voltage of the darts of color 2 starting at the 0-th edge of any copy have trivial voltage. The way to construct this graph is to apply the voltage operation in question to the symmetry type graph of a chiral $n$-polytope, which is $2^{n}_{\emptyset}$. See [8, Theorem 6.1] to see how to determine the voltage assignment.
@@
Figure 16: The voltage graph resulting from applying the operation of Section 8.1 to $2^{n}_{\emptyset}$.
@@
Let $(\mathcal{X},\Gamma(\mathcal{P}),\xi)$ be the voltage graph in Figure 16. Observing that only the edges of color two can have non-trivial voltage we can see that the facets of the derived graph will all be $n$-gons. If we remove the edges of color 2 from $\mathcal{X}$ we get two connected components. This means that $\Gamma(\mathcal{P})$ acts with two orbits on the set of facets. On the other hand, the number of connected components after removing the edges of color 0 depends on the parity of $n$: for odd $n$ we get one connected component, while for even $n$ we get two connected components.
Again, by observing that no non-trivial permutation of the set $\{\tau_{i}^{\pm 1}:1\leq i\leq n-1\}$ induces an automorphism of $\Gamma(\mathcal{P})$ we can see that no automorphism of $\mathcal{X}$ lifts, and therefore $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
@@
Notice also that, since every edge of color 2 connects the two components of $\mathcal{X}_{\overline{2}}$, we know that every 1-face (edge) of $\mathcal{X}^{\xi}$ is shared by facets in different orbits.
@@
Therefore, if $n$ is odd then $\mathcal{X}^{\xi}$ is an alternating semiregular maniplex with trivial facet stabilizer. Now we want to prove that it is also polytopal.
To do this we will use the following generalization of 6.1 (the proof is identical):
@@
Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $(r+1)$-premaniplex. Suppose $\xi(d)=1$ for every dart $d$ of color smaller than $r$. Then $\mathcal{X}^{\xi}$ is a polytope if and only if the connected components of $\mathcal{X}_{\overline{r}}$ are polytopes, and whenever there is a path $W$ with voltage 1 and colors in $[m,r]$ ending in the same connected component of $\mathcal{X}_{\overline{r}}$ as it started, there is also a path $W^{\prime}$ with the same endpoints as $W$ that uses only colors in $[m,r-1]$.
Let us denote by $\phi$ (lowercase) the right flag of the outer $n$-gon. If we start a path $W$ at $\phi$ and follow a path that alternates between colors 1 and 2 (starting with color 1), we will see that its voltage is $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$, where the subscript of $\tau_{i}$ is taken modulo $n-1$ (using $\tau_{n-1}$ when $i\equiv 0\mod n-1$), $\epsilon_{i}$ is $-1$ whenever $i\equiv n+1\mod 2n$ and 1 in any other case and $j$ is the number of edges of color 2 with non-trivial voltage used by $W$. By using 7.1 we get that the voltage of $W$ acts like a $2t$-step rotation in the Petrie poligon of $\Phi$ when $j=2(n-1)t$. Using the same argument as in 7.2 we get that if $W$ has trivial voltage, then $j$ is divisible by $2(n-1)$. This means that the last dart with non-trivial voltage used by $W$ is the dart of color 2 ending on the right flag of the $(n-1)$-th edge of the inner $n$-gon. (which is $r_{1}r_{2}\phi$). Therefore, $W$ must end in either $r_{1}r_{2}\phi$, $r_{2}\phi$, $\phi$ or $r_{1}\phi$, but only the last two are in the same connected component of $\mathcal{X}_{\overline{2}}$ as $\phi$. So, by using 9.1 we get the following result.
@@
Let $\mathcal{P}$ be a chiral $n$-polytope with simple Petrie polygons. Let $\mathcal{O}$ be the voltage operation of Section 8.1. Then $\mathcal{O}(\mathcal{P})$ is an alternating semiregular polyhedron with trivial facet stabilizer and its symmetry type graph is the one in Figure 16.
@@ 10 Higher rank
We can generalize family 1 to construct some examples of semiregular maniplexes with trivial facet stabilizer and exactly one orbit on facets.
@@
Let $\mathcal{M}$ be a regular $n$-maniplex with facets isomorphic to an $(n-1)$-maniplex $\mathcal{K}$. Suppose that $\Gamma(\mathcal{K})$ has an involution $\sigma$ in its center. Given a flag $\Phi$ in a facet $F$ of $\mathcal{M}$ we define $r_{n}\Phi$ as $\Phi\tau_{F}^{-1}\sigma\tau_{F}$, where $\tau_{F}$ is some isomorphism from $\mathcal{K}$ to $F$. One can prove that $r_{n}$ does not depend on the choice of $\tau_{F}$ (however, it does depend on the choice of $\sigma$). We may call $r_{n}\Phi$ the opposite flag of $\Phi$ in the facet $F$. Moreover, if $G$ is an $i$-face incident to $F$ with $i<n-1$, we can define the opposite face of $G$ in $F$ in a natural way. By adding edges of color $n$ between each flag and its opposite (in its respective facet), we get an $(n+1)$-premaniplex $\mathcal{X}$ (in fact, it is usually a maniplex). If $(\mathcal{X})_{\overline{0}}$ is connected, then we can use $\mathcal{X}$ to construct a semiregular maniplex with symmetry type $\mathcal{X}$:
@@
Let $F_{0},F_{1},\ldots,F_{N-1}$ be the facets of $\mathcal{M}$. Let $\mathcal{P}$ be a regular polytope of rank $N$. Let $\Gamma(\mathcal{P})=\langle\rho_{0},\ldots,\rho_{N-1}\rangle$. We construct a voltage assignment $\xi:D(\mathcal{X})\to\Gamma(\mathcal{P})$ by giving trivial voltage to all darts of color smaller than $n$, and giving voltage $\rho_{i}$ to the darts of color $n$ in the facet $F_{i}$. Then $\mathcal{X}^{\xi}$ will be an $(n+1)$-maniplex whose facets are isomorphic to $\mathcal{M}$. If no automorphism of $\mathcal{X}$ induces an automorphism of $\Gamma(\mathcal{P})$, we can be sure that $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$, immplying that $\mathcal{X}^{\xi}$ has trivial facet stabilizer. If $\mathcal{P}$ is non-degenerate, the only possible non-trivial automorphism induced by a permutation of the generators of $\mathcal{P}$ would be one that maps $\rho_{i}$ to $\rho_{N-1-i}$. In most cases, we will be able to find a labeling of the facets of $\mathcal{M}$ in a way that there is no automorphism that induces this permutation. In the rare case that this happens to be impossible, by demanding $\mathcal{P}$ to not be self-dual we make sure that this permutation of generators does not induce an automorphism of $\Gamma(\mathcal{P})$.
@@
Just as we did in Section 8.1, we may interpret this construction as applying some voltage operation that turns regular $N$-polytopes into semiregular, facet-transitive $(n+1)$-maniplexes with trivial facet stabilizer. Applying this operation to a chiral $N$-polytope would give an alternating semiregular $(n+1)$-maniplex with trivial facet stabilizer (just as in Section 9.
@@
The problem comes with polytopality. We can be certain that we need $\mathcal{M}$ to be a polytope for $\mathcal{X}^{\xi}$ to be a polytope. However it is not obvious what other conditions are necessary or sufficient to make sure that $\mathcal{X}^{\xi}$ is a polytope. This is an interesting problem to explore in the future. We have tried doing some experimentation with RAMP [19], but the examples become too big too quickly.
@@ Acknowledgments
This paper was completed while the author had a postdoctoral position at the Department of Mathematics of Northeastern University.

### v2

@@ 9 Alternating semiregular polyhedra with trivial facet stabilizer
In Section˜8.1 we have discussed how polyhedra with symmetry type graph in family 1 can be obtained by applying the voltage operation illustrated in Figure˜8 to a regular polytope of rank $n$. If instead we apply this same operation to a chiral polytope $\mathcal{P}$ with automorphism group $\Gamma(\mathcal{P})=\langle\tau_{1},\tau_{2},\ldots,\tau_{n-1}\rangle$, where once again $\Phi\tau_{i}=r_{i}r_{0}\Phi$ for all $i$ and some fixed base flag $\Phi$, then ˜8.1 tells us that the result is the derived graph of the voltage graph in Figure˜16. This graph is obtained by taking two $n$-gons and then connecting corresponding edges with a ‘‘cross’’ of edges of color 2. That is, the left flag of the $i$-th edge of one copy is 2-adjacent to the right flag of the $i$-th edge of the second copy and vice-versa. For $i>0$ the darts of color $2$ that start on the $i$-th edge of the first copy have voltage $\tau_{i}$, and the voltage of the darts of color 2 starting at the 0-th edge of any copy have trivial voltage. The way to construct this graph is to apply the voltage operation in question to the symmetry type graph of a chiral $n$-polytope, which is $2^{n}_{\emptyset}$. See ˜8.1 to see how to determine the voltage assignment.
@@
Figure 16: The voltage graph resulting from applying the operation of Section˜8.1 to $2^{n}_{\emptyset}$.
@@
Let $(\mathcal{X},\Gamma(\mathcal{P}),\xi)$ be the voltage graph in Figure˜16. Observing that only the edges of color two can have non-trivial voltage we can see that the facets of the derived graph will all be $n$-gons. If we remove the edges of color 2 from $\mathcal{X}$ we get two connected components. This means that $\Gamma(\mathcal{P})$ acts with two orbits on the set of facets. On the other hand, the number of connected components after removing the edges of color 0 depends on the parity of $n$: for odd $n$ we get one connected component, while for even $n$ we get two connected components.
@@
Notice that, since every edge of color 2 connects the two components of $\mathcal{X}_{\overline{2}}$, we know that every 1-face (edge) of $\mathcal{X}^{\xi}$ is shared by facets in different orbits.
Let $\phi$ (lowercase) be the right flag of the edge $e_{0}$ of the outer $n$-gon. We will use $\phi$ as the base flag in $\mathcal{X}$. One can find a closed path $W_{i}$ based at $\phi$ with voltage $\tau_{i}$ as follows: First go from $x$ to a flag in the edge $e_{i}$ of the outer $n$-gon by using darts of colors 0 and 1. Second, use a dart $d_{i}$ of color 2 to get to the edge $e_{i}$ of the inner $n$-gon. Third, use edges of colors 0 and 1 to get to the left flag of the edge $e_{0}$ of the inner $n$-gon. And fourth, use the dart $d_{0}$ of color 2 starting at this flag to go back to $\phi$.
Notice that if $\sigma$ is an automorphism of $\mathcal{X}$, then $\xi(W_{i}\sigma)=\xi(d_{0}\sigma)\xi(d_{i}\sigma)$. Notice also that every closed path based at $\Phi$ can be factored into paths of the form $W_{i}^{\pm 1}$. So, by knowing the voltage of the paths $W_{i}$ we can calculate the voltage of any other closed path based at $\phi$. Then, by ˜4.1, $\sigma$ lifts if and only if the mapping $\sigma^{*}:\tau_{i}\mapsto\xi(d_{0}\sigma)\xi(d_{i}\sigma)$ extends to an automorphism of $\Gamma$.
We want to make sure that if $\sigma$ is a non-trivial automorphism of $\mathcal{X}$, then it does not lift. To do this, we first notice that the involution $\tau_{n-1}$ is mapped by $\sigma^{*}$ to $\xi(d_{0}\sigma)\xi(d_{(n-1)}\sigma)$. Then $\tau_{n-1}\sigma^{*}=\tau_{i}^{\pm 1}\tau_{i\pm 1}^{\pm 1}$ for some $i\in\{0,\ldots,n-1\}$ and some choice of signs $\pm$ (subscripts are taken modulo $n$). This product is an involution if and only if one of the factors is $\tau_{0}$ (the identity) and the other one $\tau_{n-1}$ (this is again a consequence of the fact that there are no degenerate chiral polytopes). Since $\sigma$ is not the identity, this leaves us with three possibilities:
Case 1: $\sigma$ switches the inner and the outer $n$-gons (mapping $\phi$ to $\phi^{0,2}$). In this case $\tau_{i}\sigma^{*}=\tau_{i}^{-1}$. In this case, $\sigma^{*}$ cannot be extended to an automorphism of $\Gamma$.
Case 2: $\sigma$ is the reflection on a line through the left corner of $e_{0}$ (mapping $\phi$ to $\phi^{0,1,0}$). In this case $\sigma^{*}$ maps $\tau_{i}$ to $\tau_{n-1}\tau_{n-1-i}$. We can see that $\tau_{n-1}\tau_{n-1-i}$ maps $\Phi$ (the base flag of $\mathcal{P}$) to $\Phi^{0,n-1-i,n-1,0}$. This means that in the dual $\mathcal{P}^{*}$ it maps $\Phi$ to $\Phi^{n-1,i,0,n-1}$. If we now look at the action on $\Phi^{n-1,0}$ in $\mathcal{P}^{*}$ we see that $\Phi^{n-1,0}(\tau_{i}\sigma^{*})=\Phi^{n-1,i}=(\Phi^{n-1,0})^{0,i}$. In other words, $\sigma^{*}$ maps each distinguished generator of $\mathcal{P}$ to the corresponding generator of $\mathcal{P}^{*}$ when we use $\Phi^{n-1,0}$ as the base flag. Then $\sigma^{*}$ can be extended to an automorphism of $\Gamma$, if and only if $\mathcal{P}$ is properly self-dual (that is, there is an isomorphism from $\mathcal{P}$ to $\mathcal{P}^{*}$ that maps each flag to a flag in the same orbit, see [20, Section 2.2.2]).
Case 3: $\sigma$ is the composition of the previous two cases (mapping $\phi$ to $\phi^{0,1,2}$). In this case $\tau_{i}\sigma^{*}=\tau_{n-1}\tau_{n-1-i}^{-1}$, which maps $\Phi$ to $\Phi^{n-1-i,n-1}$. This means that in the dual $\mathcal{P}^{*}$ it maps $\Phi$ to $\Phi^{i,0}$. If we now look at the action on $\Phi^{0}$ in $\mathcal{P}^{*}$ (which is $\Phi^{n-1}$ in $\mathcal{P}$) we see that $\Phi^{0}(\tau_{i}\sigma^{*})=\Phi^{i}=(\Phi^{0})^{0,i}$. In other words, $\sigma^{*}$ maps each distinguished generator of $\mathcal{P}$ to the corresponding generator of $\mathcal{P}^{*}$ when we use $\Phi^{0}$ as the base flag. Then, $\sigma^{*}$ could be extended to an automorphism of $\Gamma$, if and only if $\mathcal{P}$ is improperly self-dual (that is, there is an isomorphism from $\mathcal{P}$ to $\mathcal{P}^{*}$ that maps each flag to a flag in the opposite orbit, see [20, Section 2.2.2]).
Therefore we get the following lemma:
Lemma 9.1.
The maniplex $\mathcal{X}^{\xi}$ derived from the voltage graph $(\mathcal{X},\xi)$ in Figure˜16 has symmetry type graph $\mathcal{X}$ if and only if $\mathcal{P}$ is not self-dual.
@@
In particular, if $n$ is odd then $\mathcal{X}^{\xi}$ is an alternating semiregular maniplex, and if $\mathcal{P}$ is not self-dual, then $\mathcal{X}^{\xi}$ has trivial facet stabilizer. Now we want to prove that it is also polytopal.
To do this we will use the following generalization of ˜6.1 (the proof is identical):
@@
Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $(r+1)$-premaniplex. Suppose $\xi(d)=1$ for every dart $d$ of color smaller than $r$. Then $\mathcal{X}^{\xi}$ is a polytope if and only if the connected components of $\mathcal{X}_{\overline{r}}$ are polytopes, and whenever there is a path $W$ with voltage 1 and colors in $[m,r]$ ending in the same connected component of $\mathcal{X}_{\overline{r}}$ as it started, there is also a path $W^{\prime}$ with the same endpoints as $W$ that uses only colors in [m,r1][m,r-1].
If we start a path $W$ at $\phi$ (the base flag of $\mathcal{X}$) and follow darts that alternate between colors 1 and 2 (starting with color 1), we will see that its voltage is $\prod_{i=1}^{j}\tau_{i}^{\epsilon_{i}}$, where the subscript of $\tau_{i}$ is taken modulo n1n-1 (using $\tau_{n-1}$ when $i\equiv 0\mod n-1$), $\epsilon_{i}$ is 1-1 whenever $i\equiv n+1\mod 2n$ and 1 in any other case and $j$ is the number of edges of color 2 with non-trivial voltage used by $W$. By using ˜7.1 we get that the voltage of $W$ acts like a $2t$-step rotation in the Petrie polygon of $\Phi$ when j=2(n1)tj=2(n-1)t. Using the same argument as in ˜7.2 we get that if $W$ has trivial voltage, then $j$ is divisible by 2(n1)2(n-1). This means that the last dart with non-trivial voltage used by $W$ is the dart of color 2 ending on the right flag of the (n1)(n-1)-th edge of the inner $n$-gon, which is $r_{1}r_{2}\phi$. Therefore, $W$ must end in either $r_{1}r_{2}\phi$, $r_{2}\phi$, $\phi$ or $r_{1}\phi$, but only the last two are in the same connected component of $\mathcal{X}_{\overline{2}}$ as $\phi$. So, by using ˜9.2 we get the following result.
@@
Let $n$ be an odd natural number and $\mathcal{P}$ be a non-self-dual chiral $n$-polytope with simple Petrie polygons. Let $\mathcal{O}$ be the voltage operation of Section˜8.1. Then $\mathcal{O}(\mathcal{P})$ is an alternating semiregular polyhedron with trivial facet stabilizer and its symmetry type graph is the one in Figure˜16.
@@ 10 Higher rank
In [22], given an $n$-polytope $\mathcal{P}$, the authors construct a vertex-transitive polytope $2^{\mathcal{P}}$ whose vertex-figures are isomorphic to $\mathcal{P}$. If all the facets of $\mathcal{P}$ are isomorphic to a regular (n1)(n-1)-polytope $\mathcal{K}$, then $2^{\mathcal{P}}$ is semiregular and all its facets are isomorphic to $2^{\mathcal{K}}$. If $\mathcal{P}$ has trivial automorphism group, then $2^{\mathcal{P}}$ has trivial vertex stabilizer, which implies that $2^{\mathcal{P}}$ has as many flag-orbits as $\mathcal{P}$ has flags. The authors show that for $n\geq 3$ there are arbitrarily large polytopes with trivial automorphism group and isomorphic regular facets, therefore showing that there are semiregular $(n+1)$-polytopes with arbitrarily many flag-orbits. However, $2^{\mathcal{P}}$ has also as many facet-orbits as $\mathcal{P}$ has facets.
@@
Using ˜6.3 we can construct examples of 3-polytopes with trivial facet stabilizer (therefore with as many flag orbits as flags in a facet) that are also facet-transitive. In this section we will generalize family 1 to construct some examples of semiregular maniplexes of higher rank with trivial facet stabilizer and exactly one orbit on facets.
@@
Let $\mathcal{M}$ be a regular $n$-maniplex with facets isomorphic to an (n1)(n-1)-maniplex $\mathcal{K}$. Suppose that $\Gamma(\mathcal{K})$ has an involution $\sigma$ in its center. Given a flag $\Phi$ in a facet $F$ of $\mathcal{M}$ we define $r_{n}\Phi$ as $\Phi\tau_{F}^{-1}\sigma\tau_{F}$, where $\tau_{F}$ is some isomorphism from $\mathcal{K}$ to $F$. One can prove that $r_{n}$ does not depend on the choice of $\tau_{F}$ (however, it does depend on the choice of $\sigma$). We may call $r_{n}\Phi$ the opposite flag of $\Phi$ in the facet $F$. Moreover, if $G$ is an $i$-face incident to $F$ with i<n1i<n-1, we can define the opposite face of $G$ in $F$ in a natural way. By adding edges of color $n$ between each flag and its opposite (in its respective facet), we get an $(n+1)$-premaniplex $\mathcal{X}$ (in fact, it is usually a maniplex). If $\mathcal{X}_{\overline{0}}$ is connected, then we can use $\mathcal{X}$ to construct a semiregular maniplex with symmetry type $\mathcal{X}$:
@@
Let $F_{0},F_{1},\ldots,F_{N-1}$ be the facets of $\mathcal{M}$. Let $\mathcal{P}$ be a regular polytope of rank $N$. Let $\Gamma(\mathcal{P})=\langle\rho_{0},\ldots,\rho_{N-1}\rangle$. We construct a voltage assignment $\xi:D(\mathcal{X})\to\Gamma(\mathcal{P})$ by giving trivial voltage to all darts of color smaller than $n$, and giving voltage $\rho_{i}$ to the darts of color $n$ in the facet $F_{i}$. Then $\mathcal{X}^{\xi}$ will be an $(n+1)$-maniplex whose facets are isomorphic to $\mathcal{M}$. If no automorphism of $\mathcal{X}$ induces an automorphism of $\Gamma(\mathcal{P})$, we can be sure that $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$, implying that $\mathcal{X}^{\xi}$ has trivial facet stabilizer. If $\mathcal{P}$ is non-degenerate, the only possible non-trivial automorphism induced by a permutation of the generators of $\mathcal{P}$ would be one that maps $\rho_{i}$ to $\rho_{N-1-i}$. In most cases, we will be able to find a labeling of the facets of $\mathcal{M}$ in a way that there is no automorphism that induces this permutation. In the rare case that this happens to be impossible, by demanding $\mathcal{P}$ to not be self-dual we make sure that this permutation of generators does not induce an automorphism of $\Gamma(\mathcal{P})$.
@@
Just as we did in Section˜8.1, we may interpret this construction as applying some voltage operation that turns regular $N$-polytopes into semiregular, facet-transitive $(n+1)$-maniplexes with trivial facet stabilizer. Applying this operation to a chiral $N$-polytope would give an alternating semiregular $(n+1)$-maniplex with trivial facet stabilizer (just as in Section˜9).
@@
The problem comes with polytopality. We can be certain that we need $\mathcal{M}$ to be a polytope for $\mathcal{X}^{\xi}$ to be a polytope. However it is not obvious what other conditions are necessary or sufficient to make sure that $\mathcal{X}^{\xi}$ is a polytope. This is an interesting problem to explore in the future. We have tried doing some experimentation with RAMP, but the examples become too big too quickly.
@@ Acknowledgments
This paper was completed while the author had a postdoctoral position at the Department of Mathematics of Northeastern University. The author wishes to thank Gabe Cunningham for his help with RAMP. The author also thanks the referees for their observations, in particular referee Y for helping on finding a mistake on a previous version of ˜7.2.